Simplify and express each of the following in exponential form
((5^2 )^3 × 5^4 ) ÷ 5^7

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(5^2)^3can be written as = (5)^2 x 3 ... [∵(a^m)^n = a^mn ]
= 5^6
Then,
= (5^6× 5^4) ÷ 5^7
= (5^6+4)÷ 5^7 ... [∵(a^m x a^n ) = a^m+n
= 5^10÷ 5^7
= 5^10-7 ... [∵a^m÷a^n = a^m-n]
=5^3

Thanks